OpenStudy (anonymous):

limit x-->1{ [ 1-x^1/2] / (cos^-1 x)^2 }

OpenStudy (anonymous):

Like this?\[\Huge \lim_{x \rightarrow 1}\frac{1-\sqrt{x}}{\left( \cos^{-1}x \right)^2}\]

OpenStudy (anonymous):

yes